OpenStudy (anonymous):

Identify the sequence graphed below and the average rate of change from n = 1 to n = 3. coordinate plane showing the point 2, 10, point 3, 5, point 4, 2.5, and point 5, 1.25

OpenStudy (anonymous):

an = 20(one half)n − 1; average rate of change is fifteen halves an = 10(one half)n − 1; average rate of change is fifteen halves an = 20(one half)n − 1; average rate of change is negative fifteen halves an = 10(one half)n − 1; average rate of change is negative fifteen halves

OpenStudy (anonymous):

I WIll GIVE MEDALS